OpenStudy (anonymous):

Simplify. ( - 4) (9) - (5) ( - 2)

OpenStudy (anonymous):

HELP PLS

OpenStudy (anonymous):

well what is -4 (times) 9?

OpenStudy (anonymous):

dont do this to me rosedryer

OpenStudy (anonymous):

It's a very simple question. Just use a calculator.

OpenStudy (anonymous):

its 36 leave me alone

OpenStudy (anonymous):

-36. And 5 (times) -2=-10 So -36-(-10)= -36+10= -26

OpenStudy (anonymous):

god bless